2. A 30 foot ladder weighing 100 lbs having its center of mass one-third of the way up from the bottom rests against a smooth wall so that it makes an angle of 60 degrees with the ground. If the coefficient of friction between the ground and the ladder is 0.4, how high can a 150-lb man go before the ladder slips. 


1
Expert's answer
2021-03-16T11:36:37-0400

"\vec{F}+\vec{N_0}+m\vec{g}+M\vec{g}+\vec{N_l}=\vec{0},"

"\begin{cases}
 -F+N_l=0 \\
 -mg-Mg+N_0=0
\end{cases}"

"\begin{cases}
 F=N_l\\
 (m+M)g=N_0
\end{cases}"

"F\leqslant \mu N_0"

"\begin{cases}
 N_l\leqslant \mu N_0\\
 (m+M)g=N_0
\end{cases}"

"N_l\leqslant \mu(m+M)g"


"mg\frac l3 cos \alpha +Mgxcos\alpha=N_l lsin\alpha"

"gcos\alpha (Mx+m\frac l3)=N_l lsin \alpha"

"Mx+m\frac l3=\frac{N_l ltan\alpha}{g} \leqslant\mu(m+M)l tan\alpha"

"x\leqslant l(\mu tan \alpha(1+\frac mM)-\frac{m}{3M})=27.97~ft."